# Auto-Calibration and Biconvex Compressive Sensing with Applications to Parallel MRI

### v1

@@ 6.4 Proof of Lemma (7)
6.5 Proof of Lemma (8)
We consider the case when the spanning coefficients are chosen uniformly random from a fixed basis. The structure of the proof is similar to the Gaussian case. Suppose the columns of H are chosen uniform randomly with replacement from the columns of $\textbf{W}\in\mathbb{C}^{k\times k}$ where $\textbf{W}^{*}\textbf{W}=\textbf{I}_{k}$.
@@ 6.5 Proof of Lemma (8)
Let $\textbf{A}^{\dagger}=(\textbf{A}_{S}^{*}\textbf{A}_{S})^{-1}\textbf{A}_{S}^{*}$. Suppose there exist an $\alpha$ such that $\max\limits_{j\in S^{c}}\|\textbf{A}^{\dagger}\textbf{A}_{S_{j}}\|_{F}\leq\alpha$. We postpone the estimate of the $\alpha$ to the latter part of the proof.
@@
By the Cauchy Schwartz inequality, for any y, $\|$$\tilde{X}$$\textbf{y}\|\leq\|\textbf{H}\|_{2\rightarrow 2}\|\operatorname{sgn}(\textbf{z})\|_{2}\|\textbf{y}\|_{2}\leq\sqrt{n}\|\textbf{H}\|_{2\rightarrow 2}$.
We want to bound the two terms, $\|\textbf{H}\|_{F}$ and $\|\textbf{H}\|_{2\rightarrow 2}$. For $\|\textbf{H}\|_{F}$, since each columns of H is has norm $1$, we have $\|\textbf{H}\|_{F}=\sqrt{C}$.
@@
The case when the coefficients $\textbf{h}_{l},l\in[C]$ are chosen independently and uniformly at random from the complex sphere $S_{\mathbb{C}}^{k-1}$ follows the exact argument by noticing that $\textbf{h}_{l}=\frac{\textbf{s}_{l}}{\|\textbf{s}_{l}\|_{2}}$ where $\textbf{s}_{l}$ are independent Gaussian vector drawn from $\frac{1}{\sqrt{2}}\mathcal{N}(0,1)+i\frac{1}{\sqrt{2}}\mathcal{N}(0,1)$. $\square$
@@ 7.1 Conditioning of random submatrices
The next lemma is related to the conditioning of random block submatrices for any fixed matrix $\boldsymbol{\Phi}$.
@@
[…]
@@
However, take into account the structure of our specific $\boldsymbol{\Phi}$, $\boldsymbol{\Phi}=\begin{bmatrix}\text{diag}(\boldsymbol{\psi}_{1})\textbf{B}&\cdots&\text{diag}(\boldsymbol{\psi}_{N})\textbf{B}\end{bmatrix}$. The condition on $\mu_{block}$ is strict.

### v2

@@ 6.4 Proof of Lemma 7
6.5 Proof of Lemma 8
We consider the case when the spanning coefficients are chosen uniformly at random from a fixed basis. The structure of the proof is similar to the Gaussian case. Suppose the columns of H are chosen uniformly at random with replacement from the columns of $\textbf{W}\in\mathbb{C}^{k\times k}$ where $\textbf{W}^{*}\textbf{W}=\textbf{I}_{k}$.
@@ 6.5 Proof of Lemma 8
Let $\textbf{A}^{\dagger}=(\textbf{A}_{S}^{*}\textbf{A}_{S})^{-1}\textbf{A}_{S}^{*}$. Suppose there exists an $\alpha$ such that $\max\limits_{j\in S^{c}}\|\textbf{A}^{\dagger}\textbf{A}_{S_{j}}\|_{F}\leq\alpha$. We postpone the estimate of the $\alpha$ to the latter part of the proof.
@@
By the Cauchy-Schwartz inequality, for any y, $\|$$\tilde{X}$$\textbf{y}\|\leq\|\textbf{H}\|_{2\rightarrow 2}\|\operatorname{sgn}(\textbf{z})\|_{2}\|\textbf{y}\|_{2}\leq\sqrt{n}\|\textbf{H}\|_{2\rightarrow 2}$.
We want to bound the two terms, $\|\textbf{H}\|_{F}$ and $\|\textbf{H}\|_{2\rightarrow 2}$. For $\|\textbf{H}\|_{F}$, since each column of H has norm $1$, we have $\|\textbf{H}\|_{F}=\sqrt{C}$.
@@
The case when the coefficients $\textbf{h}_{l},l\in[C]$ are chosen independently and uniformly at random from the complex sphere $S_{\mathbb{C}}^{k-1}$ follows the exact argument by noticing that $\textbf{h}_{l}=\frac{\textbf{s}_{l}}{\|\textbf{s}_{l}\|_{2}}$ where $\textbf{s}_{l}$ are independent Gaussian vectors drawn from $\frac{1}{\sqrt{2}}\mathcal{N}(0,1)+i\frac{1}{\sqrt{2}}\mathcal{N}(0,1)$. $\square$
@@ 7.1 Conditioning of random submatrices
The next lemma is related to the conditioning of random block submatrices for any fixed matrix $\boldsymbol{\Phi}$. We omit the proof in the manuscript.
@@
Suppose $\mathbf{\Phi}=\begin{bmatrix}\mathbf{\Phi}_{1}&\cdots&\mathbf{\Phi}_{N}\end{bmatrix}$, $\mathbf{\Phi}_{i}\in\mathbb{C}^{N\times k}$ and $\mathbf{\Phi}_{i}^{*}\mathbf{\Phi}_{i}=\mathbf{I}_{k}$ for all $i$. Let $S$ be a subset of $[N]$ selected at random according to the Bernoulli model with $\mathbb{E}\text{Card(S)}=n$ or uniformly random from all subset of $[N]$ of cardinality $n$. Let $P=\text{diag}[\delta_{1},\dots,\delta_{N}]$, where $\delta_{i}$ are i.i. d Bernoulli variables with $\mathbb{E}\delta_{i}=\frac{n}{N}$.
@@
Note that the derivation of the Lemma does not depend on any structures of $\Phi$. However, take into account the structure of our specific $\boldsymbol{\Phi}$, defined as $\boldsymbol{\Phi}=\begin{bmatrix}\text{diag}(\boldsymbol{\psi}_{1})\textbf{B}&\cdots&\text{diag}(\boldsymbol{\psi}_{N})\textbf{B}\end{bmatrix}$. The condition on $\mu_{block}$ is strict. Note that the derivation of the Lemma does not depend on any structures of $\Phi$. Recall that our specific $\boldsymbol{\Phi}$ is defined as $\boldsymbol{\Phi}=\begin{bmatrix}\text{diag}(\boldsymbol{\psi}_{1})\textbf{B}&\cdots&\text{diag}(\boldsymbol{\psi}_{N})\textbf{B}\end{bmatrix}$. When examining the derived conditions for our specific $\boldsymbol{\Phi}$, we find that it does not readily satisfy the necessary requirements; this makes the Lemma less useful in this context.
7.2 Algorithm
Recall from (26), $\|\mathbf{X}\|_{1,2}=\sum\limits_{j\in[N]}\|\mathbf{X}_{T_{j}\times[C]}\|_{F}$. Let $f(\mathbf{X})=\frac{1}{2}\|\textbf{Y}-\textbf{A}\textbf{X}\|_{F}^{2}$ and $g(\mathbf{X})=\lambda\|\textbf{X}\|_{1,2}$, an basic update of ISTA [Bubeck_2009] for problem (29) at iteration $k$ using stepsize $L$ is,
$\mathbf{X}_{k}=\operatornamewithlimits{argmin}_{\mathbf{X}}\{g(\mathbf{X})+\frac{L}{2}\|\mathbf{X}-[\mathbf{X}_{k-1}-\frac{1}{L}\nabla f(\mathbf{X}_{k-1})]\|_{F}^{2}\}.$ (46)
Together with $\nabla f(\mathbf{X})=\mathbf{A}^{*}(\mathbf{A}\mathbf{X}-\mathbf{Y})$ and define $\mathbf{Z}_{k-1}=\mathbf{X}_{k-1}-\frac{1}{L}\mathbf{A}^{*}(\mathbf{A}\mathbf{X}_{k-1}-\mathbf{Y})$, the update (46) becomes,
$\mathbf{X}_{k}=\operatornamewithlimits{argmin}_{\mathbf{X}}\{g(\mathbf{X})+\frac{L}{2}\|\mathbf{X}-\mathbf{Z}_{k-1}\|_{F}^{2}\}.$ (47)
Now, consider the general problem of
$\mathbf{X}=\operatornamewithlimits{argmin}_{\mathbf{X}}\{\lambda\|\mathbf{X}\|_{1,2}+\frac{L}{2}\|\mathbf{X}-\mathbf{Z}\|_{F}^{2}\}.$ (48)
Note that the optimization problem (48) is separable. In particular, if we define $\mathbf{X}=\begin{bmatrix}\mathbf{X}_{1}\\ \vdots\\ \mathbf{X}_{N}\end{bmatrix}$, $\mathbf{Z}=\begin{bmatrix}\mathbf{Z}_{1}\\ \vdots\\ \mathbf{Z}_{N}\end{bmatrix}$, $\operatornamewithlimits{argmin}\limits_{\mathbf{X}}\{\lambda\|\mathbf{X}\|_{1,2}+\frac{L}{2}\|\mathbf{X}-\mathbf{Z}\|_{F}^{2}\}=\operatornamewithlimits{argmin}\limits_{\mathbf{X}_{1},\cdots,\mathbf{X}_{N}}\{\sum\limits_{i\in[N]}\lambda\|\mathbf{X}_{i}\|_{F}+\frac{L}{2}\|\mathbf{X}_{i}-\mathbf{Z}_{i}\|_{F}^{2}\}$. Furthermore,
$\begin{split}&\operatornamewithlimits{argmin}_{\mathbf{X}_{i}}\{\sum\limits_{i\in[N]}\lambda\|\mathbf{X}_{i}\|_{F}+\frac{L}{2}\|\mathbf{X}_{i}-\mathbf{Z}_{i}\|_{F}^{2}\}\\ =&\operatornamewithlimits{argmin}_{\mathbf{X}_{i}}\{\sum\limits_{i\in[N]}\frac{\lambda}{L}\|\mathbf{X}_{i}\|_{F}+\frac{1}{2}\|\mathbf{X}_{i}-\mathbf{Z}_{i}\|_{F}^{2}\}\\ =&\text{prox}_{\frac{\lambda}{L}\|\cdot\|_{F}}(\mathbf{X}_{i})=\mathbf{Z}_{i}(1-\frac{\lambda/L}{\max(\|\mathbf{Z}_{i}\|_{F},\lambda/L)})\end{split}$ (49)
[…]